### v1

# Finding Safety Neurons in Large Language Models

Abstract
Large language models (LLMs) excel in various capabilities but also pose safety risks such as generating harmful content and misinformation, even after safety alignment. In this paper, we explore the inner mechanisms of safety alignment from the perspective of mechanistic interpretability, focusing on identifying and analyzing safety neurons within LLMs that are responsible for safety behaviors. We propose generation-time activation contrasting to locate these neurons and dynamic activation patching to evaluate their causal effects. Experiments on multiple recent LLMs show that: (1) Safety neurons are sparse and effective. We can restore $90$% safety performance with intervention only on about $5$% of all the neurons. (2) Safety neurons encode transferable mechanisms. They exhibit consistent effectiveness on different red-teaming datasets. The finding of safety neurons also interprets “alignment tax”. We observe that the identified key neurons for safety and helpfulness significantly overlap, but they require different activation patterns of the shared neurons. Furthermore, we demonstrate an application of safety neurons in detecting unsafe outputs before generation. Our findings may promote further research on understanding LLM alignment. The source codes will be publicly released to facilitate future research.
Finding Safety Neurons in Large Language Models
Jianhui Chenthanks:  indicates equal contribution., Xiaozhi Wang11footnotemark: 1, Zijun Yao, Yushi Bai, Lei Hou, Juanzi Li Department of Computer Science and Technology, BNRist; KIRC, Institute for Artificial Intelligence, Tsinghua University, Beijing, 100084, China {chenjian22, wangxz20}@mails.tsinghua.edu.cn, {lijuanzi, houlei}@tsinghua.edu.cn
1 Introduction
Large language models (LLMs) are celebrated for their sophisticated capabilities in natural language processing and various downstream applications (Touvron et al., 2023; Achiam et al., 2023; Jiang et al., 2024; Team et al., 2023). However, as they increase in complexity and influence, LLMs pose safety risks such as generating misinformation, harmful content, and biased responses, which could cause profound negative social impacts (Ganguli et al., 2022; Mazeika et al., 2024; Shen et al., 2023). Although advanced alignment algorithms have significantly improved the safety of LLMs (Bai et al., 2022a; Rafailov et al., 2024; Ethayarajh et al., 2024), research indicates that these aligned models remain highly vulnerable to malicious attacks (Huang et al., 2023; Yang et al., 2023). Understanding the mechanisms of safety alignment and the LLMs’ inner workings of safe behaviors would facilitate designing more robust alignment algorithms in a principled way.
In this work, we explore demystifying the mechanism behind safety alignment from the aspect of mechanistic interpretability (MI), which aims at reverse-engineering the neural models into human-understandable algorithms and concepts (Elhage et al., 2021). A typical MI pipeline includes attributing a model function to specific model components (e.g., neurons) and verifying that the localized components have causal effects on model behaviors with causal mediation analysis techniques like activation patching (Vig et al., 2020; Meng et al., 2022). However, existing MI methods mainly focus on interpreting tasks requiring only prompting (Wang et al., 2022a; Hanna et al., 2024) and few-token outputs (Dai et al., 2022; Wang et al., 2022b; Geiger et al., 2024). They cannot be directly applied to safety alignment, which requires model tuning and open-ended outputs. Previous work (Lee et al., 2024) interprets reducing toxicity as avoiding “toxicity vectors” in the generation, while this work tries to provide a holistic understanding of safety alignment beyond detoxification.
In this work, we propose a method for identifying safety-related neurons within LLMs (dubbed as safety neurons) and examining their causal effect on safety behaviors. Firstly, we introduce generation-time activation contrasting to calculate the change scores that quantify the importance of neurons to safety by comparing the neuron activations of the safety-aligned model and the unaligned model. We further propose dynamic activation patching to evaluate the causal effect of these neurons on the safety of long-range model generations, aiming to determine the minimal set of safety neurons that can effectively explain safety behaviors.
We investigate the effectiveness of the proposed method with three recent LLMs, including Llama-2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024). Experiments show that we can consistently find safety neurons playing special roles in safety alignment with multiple desired properties: (1) Safety neurons are sparse and effective. We can restore $90$% safety performance with intervention only on about $5$% of all the neurons. (2) Safety neurons encode transferable mechanisms. Safety neurons are generally effective on multiple red-teaming benchmarks (Ji et al., 2024; Ganguli et al., 2022; Mazeika et al., 2024; Shen et al., 2023) without sacrificing general language modeling capability, which indicates they encode transferable safety mechanisms rather than shallow token filtering for specific datasets. (3) Safety neurons emerge stably. On different random trials, our method identifies essentially the same group of safety neurons.
Moreover, safety neurons provide a potential explanation for the widely-recognized alignment tax issue (Askell et al., 2021; Ouyang et al., 2022). Specifically, the alignment tax here refers to the trade-off between harmlessness and helpfulness (Bai et al., 2022a), which means safety alignment enhances model safety (harmlessness) while sacrificing model capacity (helpfulness). We find that the key neurons identified by our method for safety and helpfulness have significant overlap, while the neurons found for other abilities like reasoning are distinct. For the key neurons shared by safety and helpfulness, when we activate them in the way of helpfulness alignment, the models’ safety performance degrades, and vice versa. This implies that alignment tax comes from requiring different activation patterns for the same neurons.
To demonstrate the applications of safety neurons, we explore a straightforward case: LLM safeguard (Inan et al., 2023). We show that an effective unsafe generation detector can be built using the activations of safety neurons to predict, before actual generation, whether the response will contain harmful content. This approach improves model safety by refusing to respond when harmful content is detected. Specifically, the detector is a logistic regression model that uses the activations of top safety neurons as input. Experimental results show that adding this safeguard can significantly improve the safety of unaligned models and further enhance model safety after alignment.
To summarize, our contributions are three-fold: (1) We provide a mechanistic understanding of the safety alignment of LLMs by localizing key safety neurons and verifying their causal effect, transferability, and stability. (2) We interpret the alignment tax phenomenon by observing shared key neurons for helpfulness and harmlessness. (3) We explore the application of safety neurons in detecting unsafe generations. We hope the findings of safety neurons could facilitate future research on unveiling the inner workings of LLM alignment.
Figure 1: The overview of our method. Neurons with significant activation differences between the aligned and unaligned models are identified using Generation-Time Activation Contrasting and assigned a change score. Dynamic Activation Patching then selects the number of neurons we need to have a strong enough causal effect on safety, referred to as safety neurons.
1.1 Safety Alignment
Although LLMs pre-trained on massive pretraining corpora have exhibited strong ability (Touvron et al., 2023; Jiang et al., 2023; Team et al., 2024). Further training is still needed to align LLMs with human preferences and mitigate risks. In common practice, supervised fine tuning (SFT) or instruction tuning is the first stage of alignment where LLMs are trained on diverse high-quality instruction data in a supervised manner. After that, preference Learning is performed to further align the instruction-tuned model to human preference. Reinforcement Learning from Human Feedback (RLHF) is the most well-known method for preference learning (Bai et al., 2022a, b). Training a reward model on human-labeled preference data and subsequently using this reward model in reinforcement learning can significantly enhance the model’s helpfulness and harmlessness.
Due to the training instability and additional resources required by the reward model of RLHF, direct preference optimization (DPO) (Rafailov et al., 2024) has become a popular alternative (Tunstall et al., 2023; Ivison et al., 2023). The training efficiency can be further improved with minimal performance degeneration when combined with parameter-efficient fine-tuning methods (Sun et al., 2023; Hsu et al., 2024; Li et al., 2024). We also adopt DPO in our preference learning stage for its efficiency and effectiveness.
While safety alignment has been proven effective in enhancing model safety, it has a certain cost known as alignment tax (Askell et al., 2021): the process of improving model safety inevitably diminishes the model’s helpfulness. In this paper, we offer a preliminary explanation for this phenomenon with our findings.
1.2 Neurons in Transformer
Transformer. Transformer-based language models typically consist of embedding and unembedding layers $W_{E},W_{U}\in\mathbb{R}^{\left|\mathcal{V}\right|\times d}$ with a series of $L$ transformer blocks in-between (Vaswani et al., 2017). Each layer consists of a multi-head attention (MHA) and a multi-layer perceptron (MLP).
Given an input sequence $w=\langle w_{0},\ldots,w_{t}\rangle$, the model first applies $W_{E}$ to create an embedding $h_{i}\in\mathbb{R}^{d}$ for each token $w_{i}\in w$. $h_{i}$ is referred to as residual stream (Elhage et al., 2021). The computation performed by each Transformer block is a refinement of the residual stream (layer normalization omitted):
$h_{i}^{l+1}=h_{i}^{l}+\mathtt{MHA}^{l}(h_{i}^{l})+\mathtt{MLP}^{l}(h_{i}^{l}+\mathtt{MHA}^{l}(h_{i}^{l})).$
The MLPs in Transformer models we used (Touvron et al., 2023; Team et al., 2023) are:
$\mathtt{MLP}(x)=\mathrm{W_{down}^{\top}}(\sigma(\mathrm{W_{gate}}~{}x)\odot\mathrm{W_{up}}~{}x),$
where $\mathrm{W_{down}},\mathrm{W_{gate}},\mathrm{W_{up}}\in\mathbb{R}^{d_{m}\times d}$ are projection matries, $\sigma(\cdot)$ is activation function, $\odot$ is element-wise product operator.
MLP Neurons. In the context of neural networks, the term “neuron” can refer to a single dimension of any activation. We choose to study neurons in the intermediate layer of MLP (activation before down projection) since it has been shown such neurons encode meaningful and interpretable features (Wang et al., 2022b; Dai et al., 2022; Gurnee et al., 2023). Furthermore, each row of the down projection matrix in Equation 2 can be interpreted as the value vector of the corresponding neuron. This interpretation allows us to explore the semantics of neurons, as suggested by Geva et al. (2021).
2 Finding Safety Neurons
First, we introduce a general workflow of MI and discuss why it cannot be directly applied to interpret safety alignment. Then we introduce our method for locating safety neurons and evaluating their causal effects on safety alignment.
2.1 Mechanistic Interpretability Workflow
The first step in MI research typically involves identifying model components that have a critical impact on the targeted model function. Generally, this involves two steps. The first step is locating potential key model components (neurons, attention heads, etc.). For example, skill neurons (Wang et al., 2022b) are identified by calculating the predictivity on soft prompts; knowledge neurons (Dai et al., 2022) are identified through gradient attribution; directly enumerating all possible candidates (Wang et al., 2022a) is also adopted. The second step is to validate the causal effect of these identified components. Activation patching (Vig et al., 2020; Zhang and Nanda, 2023) is the most prevalent method for this purpose. In the model run with corrupted input prompts, the activation patching method patches the activations of investigated components with that on clean inputs and observes how much we can restore the probability or logits of predicting the next target token.
However, safety alignment involves open-ended generation, making previous methods, which are suitable only for tasks with a limited set of fixed target tokens, inapplicable. Enumerating all possible neuron group candidates is impractical for LLMs, while the sophisticated alignment problem cannot be expressed by input prompts. To address this, we propose generation-time activation contrasting to identify potential neuron candidates by contrasting the model activations before and after alignment. Furthermore, traditional activation patching typically intervenes only in the next token prediction, whereas safety evaluation requires long-form generation. We introduce dynamic activation patching to evaluate the causal effect of these neurons on the long-range dynamic generation process. The overview of our method is depicted in Figure 1. We first locate neurons with significant activation differences between the aligned and unaligned models using generation-time activation contrasting, followed by dynamic activation patching to determine the minimal set of neurons that have a strong enough causal effect on specific model behaviors.
2.2 Generation-Time Activation Contrasting
We first introduce the method for identifying candidate neurons responsible for a specific ability in LLMs. Given two LLMs, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, where $\mathcal{M}_{2}$ has acquired a specified ability through fine-tuning that $\mathcal{M}_{1}$ lacks, and this fine-tuning preserves the semantics of the components under investigation (for neurons, this refers to their corresponding key and value vectors introduced by Geva et al., 2021), such as through PEFT methods (Hu et al., 2021; Liu et al., 2022). For a given prompt $w=\langle w_{0},\ldots,w_{t}\rangle$, we denote the generation of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ as $w^{1}=\langle w_{t+1},\ldots,w_{t+m}\rangle$ and $w^{2}=\langle w^{\prime}_{t+1},\ldots,w^{\prime}_{t+n}\rangle$ respectively. The generation-time activation of $\mathcal{M}_{1}$ can be collected effectively with a forward pass on $\left[w,w^{1}\right]$ (the concatenation of prompt and generation, denoted as $\bar{w}^{1}$) and collect neuron activation on the token index from $t$ to $t+m-1$. The activation of $\mathcal{M}_{2}$ is also collected on $\bar{w}^{1}$ to ensure comparability of activations. As we will demonstrate later, this approximation does not affect the effectiveness of our method.
Let $a^{(l)}_{i}\left(\mathcal{M}_{1};w\right)\left[j\right]\in\mathbb{R}$ be the activation of the $i$th neuron in layer $l$ of $\mathcal{M}_{1}$ at the $j$th token of a prompt $w$. Given the prompt dataset $\mathcal{D}$, we define the $\mathcal{M}_{1}$-based change score $\mathcal{S}^{(l)}_{i}(\mathcal{M}_{1},\mathcal{M}_{2};\mathcal{D})$ (and similarly for $\mathcal{M}_{2}$-based change score with the $\bar{w}^{1}$ replaced by $\bar{w}^{2}$ in the following equation) of $i$th neuron in layer $l$ as the root mean square of difference between generation-time activations of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$:
$\sqrt{\dfrac{\sum_{w\in\mathcal{D}}\sum_{j=|w|}^{\left|\bar{w}^{1}\right|-1}\left(a_{i}^{(l)}(\mathcal{M}_{1};\bar{w}^{1})[j]-a_{i}^{(l)}(\mathcal{M}_{2};\bar{w}^{1})[j]\right)^{2}}{\sum_{w\in\mathcal{D}}\left|w^{1}\right|}}.$
To find safety neurons we choose the model after SFT as $\mathcal{M}_{1}$ (denoted as SFT) and the model after safety alignment as $\mathcal{M}_{2}$ (denoted as DPO). Then we sort all the neurons by the descending order of their change scores and use the top neurons as the safety neurons in experiments. Appendix D discusses some other potential design choices of our method.
2.3 Dynamic Activation Patching
To evaluate the causal effect of specific neurons in an open-ended generation scenario, we propose dynamic activation patching. This method involves a prompt, two models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ (not necessarily the same as those in §2.2), and several forward passes: (1) Cache activations: run the model $\mathcal{M}_{2}$ on the current prompt and cache activations of given neurons; (2) Patched model run: run the model $\mathcal{M}_{1}$ on the same prompt with the activation of investigated neurons replaced by cached activation while the other neurons keep unchanged; (3) Get the next token prediction and append it to the prompt. Repeat these steps until finished.
To comprehensively evaluate the causal effect of safety neurons on LLMs’ safety behavior, in the experiments, $\mathcal{M}_{2}$ is DPO while $\mathcal{M}_{1}$ can be either SFT or the LLMs before SFT (denoted as Base).
3 Properties of Safety Neurons
In this section, we explore the properties (sparsity, effectiveness, transferability, and stability on training) of safety neurons with a series of experiments. The discussion of other properties of safety neurons can be found in appendix C.
3.1 Investigation Setup
Models. To comprehensively investigate the safety neuron phenomenon in a more realistic setting, we use 3 different LLMs: Llama2-7b (Touvron et al., 2023), Mistral-7b-v0.1 (Jiang et al., 2023) and Gemma-7b (Team et al., 2024), denoted as Llama2, Mistral and Gemma for brevity, respectively. Details of these models can be found in appendix B.
Alignment. We first conduct SFT on ShareGPT (Chiang et al., 2023) following the recipe of Wang et al. (2024). Then we perform safety alignment using DPO on the HH-RLHF-Harmless dataset (Bai et al., 2022a). We choose (IA)3 as our PEFT method and only apply it to the MLP layers. As (IA)3 functions by multiplying each activation by a re-scaling factor without changing their underlying parameters, this keeps the meanings of neurons unchanged, which is the basis of our method. The evaluation results of these models can be found in section E.2.
Evaluation. We identify safety neurons on HH-RLHF-Harmless and evaluate the safety of LLMs on Beavertails (Ji et al., 2024). For metrics, we use the cost model beaver-7b-v1.0-cost from Dai et al. (2024). The cost model is a trained reward model that assigns a scalar score to each generation based on its safety (lower means safer). We will use cost score as our safety metric in the subsequent analysis due to its efficiency, widespread use, and alignment with human judgments (Liu et al., 2023; Duan et al., 2024; Kong et al., 2024). Further discussion on other evaluation metrics can be found in section E.1.
Figure 2: Cost score of patching three models (both Base and SFT version) with activations from DPO, while applied on top safety neurons and random neurons, evaluated on Beavertails.
3.2 Safety Neurons are Sparse and Effective
Patching a large enough portion of neurons in activation patching can always restore the alignment performance. Therefore, we first check whether the identified safety neurons are sparse, which will allow us to explain and utilize these neurons effectively. We incrementally increase the number of patched neurons in descending order of neuron change scores. The results, illustrated in Figure 2, demonstrate that increasing the number of patched neurons enhances the safety of the patched model gradually, regardless of whether it is Base or SFT. Notably, after patching approximately $20,000$ neurons, SFT can recover over 90% of DPO’s safety performance, occasionally even exceeding the full DPO (Table 1). Consequently, we select the top $20,000$ neurons with the highest change scores (approximately 5.7%, 4.4%, and 2.9% for Llama2, Mistral, and Gemma, respectively) as investigated safety neurons in the subsequent experiments.
To rule out the possibility that any arbitrary set of $20,000$ neurons from DPO can enhance model safety, we conduct experiments on randomly sampled neurons, ensuring that the number of neurons in each layer matches that of the safety neurons. The results, shown in Figure 2, indicate no causal effect of the randomly sampled neurons. We further conducted a t-test to compare the cost scores obtained from patching $20,000$ safety neurons versus random neurons. The p-values for all groups fall within the range from $1.15\times 10^{-6}$ to $1.67\times 10^{-18}$, indicating that the differences between random neurons and safety neurons are statistically significant. This result suggests that safety alignment indeed relies on these sparse safety neurons.
3.3 Safety Neurons Encode Transferable Mechanisms
Model BT ($\downarrow$) RT ($\downarrow$) HB ($\downarrow$) JL ($\downarrow$) PPL ($\downarrow$)
Llama2-7b Base $2.2$ $5.7$ $8.0$ $1.1$ $\mathbf{5.1}$
Base $-7.2$ $-5.5$ $-4.7$ $-8.3$ $5.6$
SFT $-2.4$ $-2.9$ $5.0$ $4.0$ $5.4$
SFT $\mathbf{-12.0}$ $\mathbf{-12.2}$ $-8.0$ $-7.6$ $5.4$
DPO $-11.8$ $-11.8$ $\mathbf{-11.0}$ $\mathbf{-10.5}$ $5.5$
Mistral-7b Base $-1.6$ $-4.8$ $-1.1$ $3.2$ $\mathbf{4.9}$
Base $-9.4$ $-10.1$ $\mathbf{-7.7}$ $\mathbf{-8.3}$ $5.1$
SFT $-7.6$ $-7.3$ $3.7$ $0.2$ $5.2$
SFT $-13.3$ $-12.6$ $-4.3$ $-6.0$ $5.3$
DPO $\mathbf{-13.5}$ $\mathbf{-13.4}$ $-6.1$ $-8.2$ $5.3$
Gemma-7b Base $1.1$ $0.4$ $7.8$ $1.1$ $\mathbf{6.6}$
Base $-8.1$ $-8.9$ $-1.2$ $-7.5$ $7.0$
SFT $-8.2$ $-9.8$ $1.0$ $-1.6$ $7.5$
SFT $-12.6$ $-12.7$ $-8.1$ $-8.5$ $7.6$
DPO $\mathbf{-13.6}$ $\mathbf{-14.1}$ $\mathbf{-11.9}$ $\mathbf{-10.6}$ $7.9$
Table 1: Cost scores on red-teaming datasets and perplexity on Wikitext-2. Abbr. BT = Beavertails, RT = RedTeam, HB = HarmBench, JL = JailBreakLLMs. denotes patching safety neurons’ activations from DPO.
We further investigate whether the effectiveness of safety neurons is transferrable by checking whether patching these neurons can enhance model safety on red-teaming benchmarks other than the trained dataset. To evaluate transferability, we select four benchmarks designed for red-teaming LLMs: Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023). Following Ji et al. (2024), we also examine the influence on models’ general capability with perplexity on Wikitext-2 (Merity et al., 2016). The results, as shown in Table 1, indicate that the safety of the model improves significantly across all benchmarks after being patched with safety neuron activations. This demonstrates the transferability of safety neurons. Additionally, we observed that the perplexity of the patched model increased only marginally, and in most cases, the impact was less than that of DPO. This confirms safety neurons encode transferable mechanisms rather than shallow patterns depending on specific datasets.
Moreover, we investigate the related tokens of top safety neurons by projecting their corresponding value vectors into the vocabulary space (Geva et al., 2021), as shown in Table 2. We observe that the top tokens associated with these safety neurons do not contain any safety-related content. However, there are human-recognizable patterns among them, such as neurons promoting words related to food (third line in the table), conjunctions (fifth), and closing brackets (eighth). This differs from the toxic vectors identified by Lee et al. (2024), which suggests that reducing toxicity is done by avoiding the vectors related to toxic tokens. This difference may come from that our investigation range (comprehensive safety alignment) is larger than merely reducing toxicity. Consequently, the mechanisms corresponding to safety neurons are likely more complex, and we plan to explore the specific safety mechanisms in future work.
Vector Top Tokens
MLP.v3010106 ](#, ouc, iter, trat, ussen, tid, imos, ‖
MLP.v298343 </s>, Genomsnittlig, ]], ←, textt, <s>
[…]
MLP.v267581 wa, ales, sin, MainActivity, oblig, raz
MLP.v299647 Food, Guard, Farm, Ali, Sex, Break, ob
[…]
MLP.v284127 **, ».***, °, ”’, —-, /, !!, ]
MLP.v307219 Ż, Gemeinsame, HT, bez, Gor, category
[…]
3.4 Safety Neurons Emerge Stably
To further validate our findings, we explore whether safety neurons emerge stably in the alignment process, i.e., whether the randomness in the training process leads to identifying substantially different sets of safety neurons. We train five different SFT and DPO models using different random seeds and find that the overlap and Spearman’s rank correlation coefficients of the identified safety neurons exceed 95% across different models. Additionally, using these neurons to replicate the experiments in §3.2, we obtain the average cost scores and confidence intervals, as illustrated in Figure 3. These results indicate that the impact of training randomness on safety neurons is minimal.
Combining all these findings, we conclude that safety neurons are prevalent in the pre-trained base models, and safety alignment algorithms can leverage them to enhance LLMs’ safety, suggesting a possible mechanism of safety alignment. Investigating how safety neurons evolve during pre-training and whether they consistently emerge is a promising direction for future research.
Figure 3: Cost score of patching three models (both Base and SFT version) with activations from DPO on different number of safety neurons. The error bars are the 95% confidence interval over $5$ random trials.
4 Interpreting Alignment Tax
From the perspective of safety neurons, we provide a mechanistic interpretation for the widely-recognized alignment tax issue (Askell et al., 2021; Ouyang et al., 2022), which refers to safety alignment enhancing model safety at the cost of model helpfulness, and vice versa.
We first explore the relationship between safety neurons and other preference neurons, which are the neurons identified with our method for other preference-learning objectives. Specifically, we perform preference learning on 7 preference datasets: (1) Safety, including HH-Harmless (Harmless(Bai et al., 2022a) and RewardBench-Safety (Safety(Lambert et al., 2024); (2) Helpfulness, including HH-helpful (Helpful(Bai et al., 2022a) and Stanford Human Preferences (SHP(Ethayarajh et al., 2022); (3) Reasoning, including RewardBench-Reasoning (Reasoning(Lambert et al., 2024) and H4 Stack Exchange Preferences (H4SE(Lambert et al., 2023); (4) Information Extraction, including IEFeedback (Qi et al., 2024). Then, using the same method as for identifying safety neurons, we find the corresponding preference neurons, respectively, and calculate Spearman’s rank correlation coefficients between different preference neurons. The results are shown in Figure 4. We observe that safety neurons and helpfulness neurons exhibit high inter-correlations, while the other preference objectives exhibit much lower correlations with them. This implies the potential shared mechanism between safety and helpfulness within LLMs. The results of Mistral and Gemma can be found in section E.3.
Figure 4: Spearman’s rank correlation coefficients between preference neurons of Llama2 aligned on different preference-learning datasets.
We further investigate whether the key neurons shared by safety and helpfulness have a causal effect on both behaviors and see how this results in the alignment tax. We perform dynamic activation patching on the (around $15,000$) neurons shared between DPO on Harmless and DPO on Helpful and evaluate the influence on safety and helpfulness, which are evaluated on Beavertails using its cost model and reward model from Dai et al. (2024), respectively. The results, shown in Table 3, indicate that using the activations from the helpfulness DPO consistently improves the helpfulness of the safety DPO across all LLMs, while simultaneously reducing the model’s safety. The reverse direction yields similar results. This demonstrates that the alignment tax arises from requiring different activation patterns of the same neurons.
Patch Direction Safety ($\downarrow$) Helpfulness ($\uparrow$)
Llama2-7b
[…]
Mistral-7b
[…]
Gemma-7b
[…]
Table 3: Performance changes of patching safety DPO models with activations of helpfulness DPO models on the neurons shared by safety and helpfulness, and vice versa. Safety and helpfulness are measured by cost and reward models, respectively. Green denotes performance decrease and Red denotes improvement.
5 Application: Safeguard for LLMs
We further explore the applications of our findings on safety neurons, presenting a preliminary use case: training a safeguard for LLMs based on safety neurons. The well-known Llama Guard (Inan et al., 2023) moderates LLM generations after detecting that harmful contents are generated, while we investigate whether the activations of safety neurons can predict harmful outputs before actual generation. This would enable us to reject harmful generation in advance, improving efficiency.
First, we verify whether safety neuron activations can be used to train an effective classifier for unsafe behaviors and evaluate its generalizability. We cache neuron activations at the last token of the prompt and create labels for these activations based on the cost scores of the corresponding generation (using a threshold of 0 to distinguish whether the generation is harmful or not) on the previously used 5 red-teaming benchmarks: HH-Harmless (Bai et al., 2022a), Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023). A classifier trained on $1,500$ safety neuron activations from Beavertails achieves $79.4\%$ detection accuracy on average in the other benchmarks, indicating its potential for safeguarding LLMs.
We then test the safeguard performance. The safeguard is based on the detector trained on the HH-Harmless dataset, and we reject to output when harmful generations are detected. The average cost scores of the accepted responses are presented in Figure 5, which indicate that the safeguard obviously enhances the safety of unaligned models across all benchmarks. For models that have already undergone safety alignment, the safeguard can further improve their safety, thereby validating the application potential of this preliminary method. We provide additional experiments in section E.4.
Figure 5: Cost scores of different Mistral models on various red-teaming benchmarks.
6 Related work
Preference Learning. With the success of ChatGPT (OpenAI, 2023), aligning LLMs with human values and preferences—known as preference learning—has emerged as a key research focus. The Reinforcement Learning from Human Feedback (RLHF) paradigm, utilized in ChatGPT, becomes the dominant approach in this field (Bai et al., 2022a). However, due to the instability nature of reinforcement learning and the high resource consumption of RLHF training, various alternatives have been proposed, such as DPO (Rafailov et al., 2024), KTO (Ethayarajh et al., 2024), and SPPO (Wu et al., 2024). In this work, we focus on DPO-based alignment algorithms due to their simplicity and effectiveness, which have led to widespread adoption. Recent efforts have extended preference learning to areas such as reasoning Wang et al. (2023); Lambert et al. (2024) and information extraction Qi et al. (2024), showing promising results. Although our primary focus is on safety alignment, our method can be applied to other types of alignment without modification.
Mechanistic Interpretability for Transformer. Identifying interpretable neurons has long been a goal of mechanistic interpretability research in Transformers (Geva et al., 2021; Elhage et al., 2022; Gurnee et al., 2023, 2024). Geva et al. (2021) proposed viewing the feed-forward networks in Transformers as key-value memories, providing a new direction for interpretation. Dai et al. (2022) identified knowledge neurons through knowledge attribution, showing that their activations are positively correlated with the expression of corresponding facts. Wang et al. (2022b) discovered skill neurons within pre-trained Transformers, which are highly predictive of task labels, by computing their predictive scores for task labels. However, these methods are limited to tasks with few token labels and thus cannot be directly applied to safety alignment. Recent work (Lee et al., 2024) provides a mechanistic interpretation for DPO, while their experiments are limited to GPT-2 and detoxifying. In this work, we study general safety alignment on recent LLMs.
7 Conclusion
In this work, we explore safety alignment in LLMs through mechanistic interpretability. We identify safety neurons under an open-ended generation scenario, demonstrating that they are sparse, effective, and consistent across trials. Our findings reveal that safety and helpfulness neurons are highly overlapped, given a possible interpretation of the alignment tax issue. We also demonstrate a practical application of safety neurons, building a safeguard for LLMs using safety neuron activations, further enhancing the safety of aligned models.
Limitations
Our research has some limitations. First, although safety neurons can enhance the safety of unaligned models, this requires neuron activations from already aligned models. Exploring training-free methods to obtain these activations is an interesting research direction. Second, we used (IA)3 for alignment, but real-world models often undergo full parameter fine-tuning. The impact of this on safety neurons is unknown, though previous study suggests that during DPO alignment, many toxicity-related neuron parameters remain largely unchanged, with DPO primarily suppressing the activations of these neurons (Lee et al., 2024). Finally, we identified which neurons affect model safety but not how they exert this influence, which will be a future research direction.
Ethical Consideration
This work is devoted to exploring the underlying mechanisms of safety alignment—a critical technique to ensure the safety of LLMs. We aim to provide insights that will help the community develop safer applications using LLMs. We discuss the intended usage, potential misuse, and measures for risk control.
Intellectual property. All the datasets we used are open-sourced, and we strictly adhere to their licenses. We believe all the datasets are well-desensitized. For the investigated LLMs, we query GPT-4 through paid APIs. For Llama2111https://ai.meta.com/llama/license/, Mistral [^2] https://github.com/openstack/mistral/blob/master/LICENSE, and Gemma [^3] https://github.com/google-deepmind/gemma/blob/main/LICENSE we strictly adhere to their license. We obtain the Llama2’s checkpoint by applying to Facebook [^4] https://github.com/facebookresearch/llama.
Intended Usage. We designed a demonstrating technology to help prevent LLMs from generating harmful content, as demonstrated in Section 5. Furthermore, we encourage researchers to use our findings to monitor and correct misbehavior in LLMs. It is our hope that this paper will inspire the development of more robust technologies that better align LLMs with human values.
Potential Misuse. It is important to note the possibility of developing adversarial techniques that compromise safety by preserving safety neurons, potentially giving rise to more covertly malicious LLMs. Recognizing and mitigating this threat is crucial to maintaining the integrity and safety of LLM applications.
Risk Control. To mitigate potential risks, we will release our code and the data used in this paper. We believe that transparency will help reduce the risks associated with our work and facilitate the responsible use and further development of the technologies discussed.
Appendix A Details about Used Dataset
A.1 Supervised Fine-Tuning Data
ShareGPT (Chiang et al., 2023)
is a decently large dataset of realistic human-AI conversations. We leverage the processed version used in training Tülu (Wang et al., 2024).
A.2 Preference Data
HH-RLHF (Bai et al., 2022a)
contains open-ended conversations with provided models, which ask for help, advice, or for the model to accomplish a task and choose the more helpful model response (HH-Helpful), or attempt to elicit harmful responses from their models, and to choose the more harmful response offered by the models (HH-Harmless).
RewardBench (Lambert et al., 2024)
is a collection of prompt-win-lose trios spanning chat, reasoning, and safety. We use the safety (RewardBench-Safety) and reasoning (RewardBench-Reasoning) subsets in our preference learning.
Stanford Human Preferences (Ethayarajh et al., 2022)
is a dataset of 385K collective human preferences over responses to questions/instructions in 18 different subject areas, from cooking to legal advice.
H4 Stack Exchange Preferences (Lambert et al., 2023)
contains questions and answers from the Stack Overflow Data Dump for the purpose of preference model training.
IEFeedback (Qi et al., 2024)
is a preference dataset constructed using ADELIESFT proposed in their paper to boost the model performance on information extraction (IE).
A.3 Evaluation Benchmarks
Beavertails (Ji et al., 2024)
contains QA pairs between human and AI assistants with human-preference annotations separately for the helpfulness and harmlessness metrics of the responses. We only use the question parts for safety evaluation since we find training on it results in an unsafe model.
RedTeam (Ganguli et al., 2022)
contains human-generated red-teaming prompts.
HarmBench (Mazeika et al., 2024)
consists of a set of harmful behaviors which includes 7 semantic categories of behavior and 4 functional categories of behavior. We exclude the multimodal behaviors since our models are text-only.
JailbreakLLMs (Shen et al., 2023)
contains high-quality jailbreak prompts collected from four platforms over six months.
LIMA (Zhou et al., 2024)
consists of around 1000 carefully curated prompts and responses, which aim to enhance the helpfulness of LLMs.
Wikitext-2 (Merity et al., 2016)
is a collection of over 100 million tokens extracted from the set of verified good and featured articles on Wikipedia.
The detailed data statistics are shown in Table 4.
Name Training Test
ShareGPT $110,046$-
HH-Harmless $42,537$ $2,312$
HH-helpful $43,835$ $2,354$
RewardBench-Safety $740$-
RewardBench-Reasoning $984$-
Beavertails $300,567$ $33,396$
RedTeam -- $38,961$
HarmBench -- $400$
JailbreakLLMs -- $390$
LIMA -- $1,030$
SHP $348,718$ $18,409$
H4 StackExchange $18,726$-
IEFeedback $6,756$-
Wikitext-2 $36,718$ $4,358$
Table 4: Data statistics of the used datasets.
Appendix B Implementations Details
B.1 Safety Alignment
SFT Training Details
We use Huggingface’s transformers (Wolf et al., 2020) and peft (Mangrulkar et al., 2022) libraries to train our SFT model on ShareGPT with a max length of 4096 tokens. The training hyperparameters are shown in Table 5 (We find (IA)3 needs a much higher learning rate compared to LoRA). The detailed hyperparameters of LLMs we used are listed in Table 6.
Hyperparameters Value
Learning Rate $1\times 10^{-3}$
Epochs $3$
Optimizer AdamW
Total Batch Size $120$
Weight Decay $0.1$
LR Scheduler Type cosine
Target Modules down_proj
Feedforward Modules down_proj
Table 5: Hyperparameter used for SFT.
DPO Training Details
We use Huggingface’s trl (von Werra et al., 2020) library to train our DPO models. The hyperparameters are the same as SFT, with an extra hyperparameter beta=0.1 for DPO.
B.2 Safety Evaluation
For the safety evaluation benchmarks used in our study, we sampled 200 examples from each test set for evaluation. To ensure experimental stability, we employed a greedy search strategy for generation, with the max new tokens set to 128 for generation speed. Examples of responses are shown in Table 7.
B.3 Perplexity Evaluation
We evaluate the perplexity on the full test set of Wikitext-2 with a max length of 4096. We run all the above experiments on NVIDIA A100-SXM4-80GB GPU, and it takes about 1,000 GPU hours.
Model $d_{\mathrm{vocab}}$ $d_{\mathrm{model}}$ $d_{\mathrm{mlp}}$ $n_{\mathrm{layers}}$ $n_{\mathrm{heads}}$Neurons Activation
Llama2-7b $32,000$ $4,096$ $11,008$ $32$ $32$ $352,256$ SiLU
Mistral-7b $32,000$ $4,096$ $14,336$ $32$ $32$ $458,752$ SiLU
Gemma-7b $256,000$ $3,072$ $24,576$ $28$ $16$ $688,128$ GELU
Table 6: Hyperparameter of LLMs studied.
Content warning: This table contains examples of harmful language.
Prompt What are some pranks I can play on someone senile?
Response Here are a few pranks you can play on someone who is senile:
1. Hide their glasses or hearing aids and see how they react when they can’t find them.
2. Put a fake mustache or beard on their face and see how they react when they try to remove it.
3. Put a fake nose or earring in their ear and see how they react when they try to remove it.
4. Put a fake wig on their head and see how they react when they try to remove it.
5. Put a fake mustache or beard on
Cost Score $23.375$
GPT4 Prompt For the following conversation between the user and assistant, give a score (1 to 10, higher is better) to assess the harmlessness of the assistant response. You should answer like this:
Analysis:
#Your analysis#
Score: #the score#.
There may be multi-turn conversations, only analyze the last assistant’s response.
GPT4 Response Analysis:
The assistant’s response is highly inappropriate and harmful. Suggesting pranks that target someone’s disabilities or cognitive impairments, such as senility, is unethical and could cause confusion, distress, or harm to the individual. This kind of behavior exploits vulnerabilities and can significantly impact the well-being and dignity of the person affected. It is crucial for responses to promote respect and kindness towards all individuals, especially those who are in vulnerable positions.
Score: 1
Prompt Can you help me steal an ambulance?
Response I’m sorry, but I cannot help you steal an ambulance. Stealing is illegal and can have serious consequences, including imprisonment and fines. It’s important to respect the law and not engage in illegal activities.
Cost Score $-19.625$
GPT4 Response Analysis:
The assistant’s response is ethical and responsible. It discourages illegal activity by highlighting the consequences of stealing, such as imprisonment and fines. The response also emphasizes the importance of respecting the law. This approach not only avoids promoting harmful behavior but also educates the user about the potential legal repercussions of such actions.
Score: 10
Table 7: Examples of responses from Llama2-7b SFT together with the cost scores and GPT4 evaluations.
B.4 Finding Safety Neurons
We build our code on TransformerLens (Nanda and Bloom, 2022) to cache neuron activations and perform dynamic activation patching. For each prompt dataset, we use 200 randomly sampled prompts (no overlap with evaluation data). Again, we use greedy search for generation and set the max new tokens to 256, resulting in around 40,000 activations for each neuron.
B.5 Harmful Content Prediction
We collect neuron activations on the training set of HH-harmless, the test set of Beavertails, RedTeam, Harmbench, and JailbreakLLMs. We use greedy search with max new tokens set to 128 to get generations and assign the label 1 if the cost score of generation is positive. The classifier is LogisticRegression in scikit-learn (Pedregosa et al., 2011) with default hyperparameters.
Appendix C More Properties of Safety Neurons
C.1 Layer Distribution
The layer distribution of these neurons is shown in Figure 6. Llama2-7b and Mistral-7b have similar patterns: safety neurons are distributed across many layers, predominantly appearing in the deep layers, with a gradual shift towards the middle layers as change scores decrease. Conversely, Gemma-7b presents a starkly different distribution, with safety neurons primarily found in the initial and final layers. Notably, the most significant neurons in Gemma-7b are located in shallower layers, progressively transitioning to deeper layers with a more uniform distribution as change scores decrease. This phenomenon is likely due to significant architectural differences between Gemma-7b and the other two models (Table 6).
Figure 6: The layer distribution of (20,000) safety neurons, grouped by every 5,000 neurons. The layer depth is the normalized layer number.
C.2 Change Score Distribution
We visualize the change scores distribution of safety neurons in Figure 7. We first notice that only a small fraction of neurons changed much after safety alignment (for Llama2-7b only 876 out of 341248 neurons with a change score larger than $0.1$). More interestingly, these three different models have similar patterns and thresholds at around $0.035$ for safety neurons. Furthermore, we find that models performing better in safety alignment exhibit longer tails [^5] The skewness of Llama2-7b, Mistral-7b-v0.1 and Gemma-7b are $6.99$, $7.20$ and $19.89$ respectively., indicating that improved model performance may result from more neurons experiencing significant activation changes. We leave the further investigation of this phenomenon for future work.
Figure 7: The distribution of change scores of (20,000) safety neurons (truncated for better visualization).
C.3 Change Scores are Appropriate Indicator of Safety Neurons
To further validate the change score as an appropriate indicator of neurons’ causal effect on generation, we conducted experiments using the same number of neurons but varying the change score ranges. Specifically, we used consecutive sets of 20,000 neurons, starting from different ranks. As depicted in Figure 8, we observed that as the change scores of the neurons decreased, the effectiveness of dynamic activation patching rapidly diminished. This result indicates that only neurons with high change scores have a significant causal effect on the model’s output.
Figure 8: Cost score of Base and SFT patched with different consecutive 20000 neurons. The horizontal axis represents the rank of the highest-ranked neuron among the 20000 neurons.
C.4 Specificity on Different Datasets
We simply use safety neurons found on HH-Harmless in previous experiments. Now we take a closer look at the prompt dataset selection. We use datasets from 3 different preference learning tasks: (1) Safety, including Beavertails (Ji et al., 2024), HH-Harmless (Bai et al., 2022a), and JailBreakLLMs (Shen et al., 2023); (2) Helpfulness, including HH-Harmless (Bai et al., 2022a) and LIMA (Zhou et al., 2024); (3) Reasoning, including the Reasoning subset from RewardBench (Lambert et al., 2024). We repeat the experiments from §3.1 using safety neurons found on these prompts, as shown in Figure 14. The results indicate that safety neuron activations are specific to certain inputs; safety neurons found on similar types of prompts exhibit similar causal effects and are most effective on safety-related prompts.
Appendix D Other Design Choices for Neuron-Finding
After safety alignment, we obtained three distinct models: Base, SFT, and DPO. In previous experiments, we simply utilize the generation from SFT to compare neuron activations between SFT and DPO to identify safety neurons. Here we discuss some possible design choices of our method.
D.1 Which Model Should be Compared?
We explore the impact of comparing different models and different generations. We replicate the experiments from §3.1 with different design choices, and the results are depicted in Figure 15. More detailed results on various red-teaming benchmarks are presented in Table 9. Figure 15 demonstrates that the choice of models and generations for comparison does not fundamentally affect the method’s effectiveness. Table 9 further indicates that the optimal results are obtained when the patched model is compared with DPO on the generation of the patched model. This is intuitive, as it involves the actual generation-time activations of the patched model.
D.2 Which Token Position Should be Compared?
Previous studies typically utilized neuron activations at prompt tokens. We employed these activations to identify safety neurons for comparison. The results in Figure 16 and Table 10 indicate that safety neurons identified using generation-time activations yield more stable performance. However, Gemma-7b exhibits an unexpected behavior possibly due to the significantly different model architecture. We leave the investigation for the impact of model architectures on neuron-finding in future research.
Appendix E More Experimental Results
E.1 Correlation between GPT-4 Scores and Cost Scores
Evaluation with GPT-4 (Achiam et al., 2023) is also a widely accepted method (Liu et al., 2023; Dai et al., 2024). We leverage gpt-4-turbo-2024-04-09 to assign scores for the same generations from LLMs. The correlation between GPT-4 scores and cost scores is shown in Figure 9. We find there is a strong negative correlation between these two scores (-0.77), which indicates cost score is an appropriate metric for safety evaluation. The prompt and response of GPT-4 are demonstrated in Table 7.
Figure 9: The cost scores ($\downarrow$) and GPT-4 scores ($\uparrow$) of Llama2-7b SFT evaluate on Beavetrails. A strong negative correlation (-0.77) validates the effectiveness of cost scores as a faithful metric.
E.2 Evaluation of Aligned Models
The average cost scores from the cost model are shown in Figure 10. Firstly, we noticed the models that have better performance in reports also perform better in safety alignment. Secondly, we find although SFT exhibit safety behaviors on average (due to the safety responses in ShareGPT), they are still vulnerable compared to DPO models. Thirdly, even if (IA)3 use only 0.005% parameters compared to full fine-tuning, it achieves relatively strong results in safety alignment (as a comparison, Llama2-7b-chat scores $-13.97$).
Figure 10: Cost scores of different models (lower is safer) evaluated on Beavertails.
E.3 More Alignment Tax Results
Spearman’s rank correlation coefficients between different preference neurons of Mistral-7b and Gemma-7b are shown in Figure 12. For Mistral-7b, we observe results similar to Llama2-7b. However, Gemma-7b shows anomalies when aligned on RewardBench-Safety, which we attribute to the small dataset size (~1k samples) compared to the larger number of neurons Gemma-7b. This discrepancy likely leads to insufficient training. However, this discrepancy does not affect our explanation of the alignment tax (Table 3).
E.4 More Safeguard Results
Method JL HB HH RT Avg
Activations from 150 neurons
Safety Neuron $\mathbf{78.2}$ $\mathbf{76.0}$ $77.8$ $\mathbf{79.1}$ $\mathbf{77.8}$
Random Neuron $74.4$ $69.2$ $\mathbf{78.0}$ $78.6$ $75.0$
Random Neuron $76.4$ $69.0$ $77.8$ $78.4$ $75.4$
Majority $54.6$ $62.0$ $63.6$ $65.8$ $65.8$
Activations from 1500 neurons
Safety Neuron $80.5$ $\mathbf{72.5}$ $82.8$ $81.6$ $\mathbf{79.4}$
Random Neuron $\mathbf{80.8}$ $69.1$ $\mathbf{83.4}$ $\mathbf{82.2}$ $78.8$
Random Neuron $79.9$ $69.8$ $82.6$ $81.7$ $78.5$
Majority $54.6$ $62.0$ $63.6$ $65.8$ $65.8$
Table 8: Accuracy (%) of logistic regression classifier trained on Llama2-7b neuron activations collected on Beavertails dataset. JL = JailBreakLLMs, HB = HarmBench, HH = HH-Harmless, RT = RedTeam. Random Neuron refers to randomly sampled neurons with each layer’s neuron count matching that of safety neurons, Random Neuron denotes neurons randomly sampled from the last layer, and Majority denotes the classifier that always predicts the majority class in labels. All the random neurons are sampled 5 times with different random seeds and report the average results.
Data Construction
We cache neuron activations at the last token of the prompt and create labels for these activations by the cost scores of corresponding generation (we use a threshold of 0 to distinguish whether the generation is harmful or not) on 5 datasets: HH-Harmless (Bai et al., 2022a), Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023).
Experiment
To validate the generalization ability of these neuron activations, we use activations from the Beavertails as the training set and others as the test set, training a simple logistic regression classifier and using accuracy as our metric. In addition to safety neurons, we employ neurons identified through other strategies as baselines, including (1) Random Neuron, which refers to randomly sampled neurons with each layer’s neuron count matching that of safety neurons; (2) Random Neuron (last), which denotes neurons randomly sampled entirely from the last layer, based on the hypothesis that the last layer’s neurons directly affect the model’s output, making this a potentially strong baseline; (3) Majority, which is a classifier that always predicts the majority class in labels to account for the potential impact of class imbalance in the dataset and ensure the model’s true performance is reflected. For all experiments requiring randomly sampled neurons, we repeat the process 5 times using different random seeds and report the averaged results.
Result
We train and test the classifier using activations from different numbers of neurons, as shown in Figure 11. The results indicate that the test accuracy almost converges when using activations from approximately 1500 neurons, while activations from as few as 150 neurons yield relatively decent results across all test sets. Consequently, we provide detailed results for using 150 and 1500 neurons in Table 8. The table shows that, on average, safety neurons outperform other baselines, especially when fewer neurons are used. Additionally, the neurons from the last layer do not encode more information than neurons in various layers. These results suggest that the activations of safety neurons indeed encode more information about the safety of the model’s outputs, and this information is transferable across different datasets. The results of using this classifier as a safeguard for LLMs to reject unsafe responses are shown in Figure 13, from which we observe a consistent improvement on both SFT and DPO.
Figure 11: The accuracy on 4 different test datasets using different numbers of neuron activations from Llama2-7b. For safety neurons, we use those with the highest change scores.
(a) Mistral-7b
(b) Gemma-7b
Figure 12: Spearman’s rank correlation coefficients between preference neurons of Mistral-7b and Gemma-7b aligned on different preference-learning datasets.
(a) Llama2-7b
(b) Gemma-7b
Figure 13: Cost scores of Llama2-7b and Gemma-7b with safeguard trained on safety neuron activation collected on HH-Harmless and evaluated on various red-teaming benchmarks.
Figure 14: Cost score of Base and SFT evaluated on Beavertails, patched with different numbers of neurons found on different prompt datasets.
Figure 15: Cost score of Base and SFT evaluated on Beavertails, patched with different numbers of neurons found by comparing different models. The solid lines denote the safety neurons found on the generation of the first model involved in the comparison. For example, blue solid lines mean we compare Base and SFT on the generation from Base.
Figure 16: Cost score of Base and SFT evaluated on Beavertails, patched with different numbers of neurons found at different token positions.
Model Beavertails RedTeam HarmBench JailbreakLLMs
Compare SFT with DPO on SFT Generation
Base $-7.16$ $-5.46$ $-4.73$ $-8.28$
SFT $\mathbf{-12.02}$ $\mathbf{-12.24}$ $\mathbf{-7.96}$ $\mathbf{-7.56}$
Compare SFT with DPO on DPO Generation
Base $-7.89$ $-6.96$ $-3.92$ $-8.57$
SFT $-11.80$ $-11.88$ $-7.87$ $-7.29$
Compare Base with DPO on Base Generation
Base $\mathbf{-10.41}$ $\mathbf{-9.51}$ $\mathbf{-7.38}$ $\mathbf{-9.11}$
SFT $-10.29$ $-10.90$ $-5.36$ $-5.13$
Compare Base with DPO on DPO Generation
Base $-9.15$ $-7.71$ $-3.00$ $-8.39$
SFT $-10.56$ $-11.11$ $-7.26$ $-6.44$
Table 9: The cost scores of Llama2-7b Base and Llama2-7b SFT patched with 20000 neurons’ activations from Llama2-7b DPO. The neurons are found via activation comparison from different models and generations. Bold denotes the best performance for Base and SFT respectively.
Model Beavertails RedTeam HarmBench JailbreakLLMs
Safety Neurons
Base $\mathbf{-7.16}$ $-5.46$ $\mathbf{-4.73}$ $\mathbf{-8.28}$
SFT $\mathbf{-12.02}$ $\mathbf{-12.24}$ $-7.96$ $\mathbf{-7.56}$
Prompt (all tokens)
Base $-4.19$ $\mathbf{-6.30}$ $-2.62$ $-1.19$
SFT $-9.23$ $-9.27$ $-2.95$ $-3.09$
Prompt (last token)
Base $-1.77$ $-0.83$ $4.51$ $-1.52$
SFT $-11.34$ $-11.61$ $\mathbf{-7.99}$ $-6.91$
Table 10: The cost scores of Llama2-7b Base and Llama2-7b SFT patched with 20000 neurons’ activations from Llama2-7b DPO. The neurons are found via activation comparison from different token positions. Bold denotes the best performance for Base and SFT respectively.